OpenStudy (anonymous):

Evaluate the line integral directly \[\oint_c xydx+x^2dy\]

OpenStudy (anonymous):

What is c?

OpenStudy (anonymous):

from 0 to 1

OpenStudy (anonymous):

(0,0) to (1,0)?

OpenStudy (anonymous):

Sorry it's (0,0) (0,1) (3,0) (3,1)

OpenStudy (anonymous):

\[\int_0^1 \int_0^3 xydx+x^2 dy?\]

OpenStudy (zarkon):

you need to parametrize your path

OpenStudy (anonymous):

what do you mean?

OpenStudy (zarkon):

have you ever done any of these problems before?

OpenStudy (anonymous):

No, I have evaluated the integral using greens theorem, but I am asked to evaluate it directly. How is that different from greens theorem?

OpenStudy (anonymous):

do I use \[\int_c f(x,y)ds\] ?

OpenStudy (anonymous):

from (0,0) to (0,1) the integral is zero. x=0 dx =0 y=t dy =dt The quantity to be integrated is zero. So the integral is zero.

OpenStudy (anonymous):

From (3,0) to (3,1) x=3 , dx =0 y=t dy =dt \[ xydx+x^2dy= 9 dt\\ \int_0^1 9 dt =9 \]

OpenStudy (anonymous):

Is your c the boundary of the rectangle?

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

From (0,0) to (3,0) y=0 dy=0 x=t dx =dt xydx+x^2dy=0 So the integral is zero

OpenStudy (anonymous):

From (3,1) to (0,1) y=1 dy =0 x=3 dx =0 xydx+x^2dy=0 The integral is zero

OpenStudy (anonymous):

but using greens theorem I get 4.5

OpenStudy (anonymous):

I also got 4.5 using Green's Theorem. Let me double check.

OpenStudy (anonymous):

Here is my mistake From (3,1) to (0,1) y=1 dy =0 x=t dx =dt xydx+x^2dy=t dt The integral \[ \int_3^0 t dt=-\frac 92 \] 0 + 9-9/2 + 0 =9/2

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

Yep sorry I had to go for a little bit, but I'm back. Oh that makes sense now.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yw